# Randomized Black-Box PIT for Small Depth +-Regular Non-commutative Circuits

### v1

@@ 4.1.2 Non-zeroness of $\hat{f}$
Let $Z=\{z_{1},\ldots,z_{n}\}$ be the set of new commuting variables. Let $g\in\mathbb{F}\mathopen{\langle}X\mathclose{\rangle}$ be a polynomial of degree $D$ computed by a $\Sigma\Pi^{*}\Sigma$ circuit of size $s$. Then $g$ can be expressed as $g=\sum_{i\in[s]}\prod_{j\in[D]}L_{ij}$, where $L_{ij}$ are homogeneous linear forms. Let $P_{i}=\prod_{j\in[D]}L_{ij}$, $i\in[s]$. We have $g=\sum_{i\in[s]}P_{i}$. For $I\subseteq[D]$ with size at most $s-1$, define $P_{i,I}$ as the polynomial obtained from $P_{i}$ by treating linear forms indexed by $I$ as noncommuting and the rest of the linear forms as commuting. We replace $x_{i}$ variable appearing in $[D]\setminus I$ by a new commuting variable $z_{i}$.
The number of n.c. variables appearing in $P_{i,I}\in\mathbb{F}[Z]\mathopen{\langle}X\mathclose{\rangle}$ is bounded by $|I|<s$. We refer to this as the n.c. degree of the polynomial $P_{i,I}$. Since this degree is small, $P_{i,I}$ can be converted into a commutative polynomial while preserving its non-zeroness. Let $P^{(c)}_{i,I}$ denote the commutative polynomial obtained from $P_{i,I}$ and define $g_{I}=\sum_{i\in[s]}P^{(c)}_{i,I}$. To keep all guesses of the set $I$ distinct, additional commutative variables $\xi=\{\xi_{1},\xi_{2},\cdots,\xi_{k+1}\}$ are introduced in [AJMR19]. The transformed commutative polynomial obtained in [AJMR19] is given by:
@@
Let $f=\sum_{i\in[s]}\prod_{j\in[D_{2}]}Q_{ij}$ be a n.c. polynomial over $X=\{x_{1},\cdots,x_{n}\}$, computed by a $\Sigma\Pi^{*}\Sigma\Pi^{*}\Sigma$ circuit of size $s$. Let $D_{1}$ denote the degree of the polynomial $Q_{ij}$, $i\in[s],j\in[D_{2}]$. Let $\hat{f}\in\mathbb{F}[Y\sqcup Z]\mathopen{\langle}\xi\mathclose{\rangle}$ be defined as $\hat{f}=\sum_{i\in[s]}\prod_{j\in[D_{2}]}\hat{Q}_{ij},$ where $\hat{Q}_{ij}$ are as defined earlier. Then, if $f\neq 0$ then $\hat{f}\neq 0$.
This above lemma 4.1.1 can be generalized. We will utilize this generalization to transform polynomials computed by larger depth $+$-regular circuits. We note this as a remark.
Remark 4.1.3.
The proof of Lemma 4.1.1 is based on the observation that if each polynomial $Q_{ij}$ (where $i\in[s]$ and $j\in[D_{2}]$, computable by a depth-3 circuit) can be converted into a n.c. polynomial $\hat{Q}_{ij}$ while preserving non-zeroness, then this lemma asserts that this transformation can be applied to each $Q_{ij}$ to create the n.c. polynomial $\hat{f}$, maintaining the non-zeroness of the entire polynomial. This can be extended to higher depths as follows: suppose $f$ is computed by the sum of the products of $Q_{ij}$’s, where each polynomial $Q_{ij}$ is computed by a depth $d$ $+$-regular circuit and each $Q_{ij}$ can be converted into a n.c. polynomial $\hat{Q}_{ij}$ while preserving non-zeroness, then each $Q_{ij}$ can be converted to $\hat{Q}_{ij}$ to create the n.c. polynomial $\hat{f}$, maintaining the non-zeroness of the entire polynomial. The proof follows a similar structure to that of Lemma 4.1.1
The resulting n.c. polynomial $\hat{f}$ still has an exponential degree in $\xi$ variables, but each $\hat{Q}_{ij}$ is structured as $s$-ordered power-sum polynomials. Importantly, $\hat{f}$ does not contain any monomials from the spurious polynomial $F=\sum_{\begin{subarray}{c}m\in Mon(f)\end{subarray}}F_{m}$.
@@ 4.1.3 Small degree case
When $D_{1}<s-1$, we can use a small substitution automaton of size $c$, bounded by $s$, to substitute fresh n.c. double-indexed variables at each position within each $Q_{ij}$. Let $Z=\{z_{\ell k}\mid\ell\in[c]\text{ and }k\in[n]\}$ be the new set of n.c. variables. The automaton replaces the variable $x_{k}$ in the $\ell$-th position of a monomial of $Q_{ij}$ with $z_{\ell k}$. The resulting polynomial is denoted by $\hat{Q}_{ij}$, which remains n.c. over $Z$. Unlike the high-degree case, in the small-degree case, the automaton can identify the boundary between $Q_{ij}$ and $Q_{i,j+1}$, ensuring that no spurious monomials are produced, that is, $F=0$.
@@
However, it is important to note that this transformation alone will not provide a black-box PIT, as we cannot guarantee the non-zeroness of the sum of products of these $\hat{Q}_{ij}$ polynomials if we simply treat all $\hat{Q}_{ij}$ as commutative. At this stage, the variables in $Z$ are still considered n.c. in the polynomial $\hat{f}$.
@@ 4.2 Step 2: Product Sparsification
In the second step of our transformation, we focus on the sparsification of the n.c. polynomial $\tilde{f}\in\mathbb{F}\mathopen{\langle}\xi\mathclose{\rangle}$ with respect to the $s$-ordered power-sum polynomials $\hat{Q}_{ij}$. This transformation affects both the good part $\hat{f}_{1}$ and the spurious part $F_{1}$ of the polynomial $\tilde{f}$.
We begin by analyzing the transformation of $\hat{f}_{1}$, which is defined as:
@@
where each $\hat{Q}_{ij}$ is an $s$-ordered power-sum polynomial in the n.c. variables $\xi=\{\xi_{1},\ldots,\xi_{s}\}$.
The key observation is that we can preserve the non-zeroness of $\hat{f}_{1}$ by retaining at most $s-1$ of the $s$-ordered power-sum polynomials $\hat{Q}_{ij}$ in each product $\prod_{j\in[D_{2}]}\hat{Q}_{ij}$ as n.c. while treating the remaining ones as commutative. This is stated in the following lemma, which we refer to as the product sparsification lemma. This lemma generalizes Lemma 6.2 from [AJMR19]. Unlike [AJMR19], we are dealing with the product of n.c. polynomials where the degree of individual factors can be greater than 1 if we consider them as commutative they may become 0. The proof of this lemma is crucially used the Claim 4.0.1, and the proof can be found in Appendix B. We will revisit the combination of $\hat{f}_{1}$ and $F_{1}$ in LABEL:{step1.1} to complete our analysis.
@@
where each $\hat{Q}_{ij}$ is an $s$-ordered power-sum polynomial over $\xi=\{\xi_{1},\xi_{2},\ldots,\xi_{s}\}$. Then, there exists a subset $I\subseteq[D_{2}]$ with size at most $s-1$ such that if we treat the polynomials $\hat{Q}_{ij}$ for $j\in I$, as non-commutative and the others($j\not\in I$) as commutative, then the polynomial $\hat{f}_{1}$ remains non-zero.
@@
The proof of Lemma 4.2.1 relies only on the fact that each $\hat{Q}_{ij}$ where $i\in[s]$ and $j\in[D_{2}]$, is an $s$-ordered power-sum polynomial. The size of the index set $I$ depends only on the number of summands $s$. Importantly, the proof does not depend on the fact that each $\hat{Q}_{ij}$ is obtained from a polynomial computed by a $\Sigma\Pi^{*}\Sigma$ circuit. This allows us to generalize the lemma to sums of products of arbitrary $k$-ordered power-sum polynomials where $k\in\mathbb{N}$, with a proof that is analogous to that of the original lemma. We will utilize this generalization to transform polynomials computed by higher-depth $+$-regular circuits.
Since we do not know the index set $I$, the substitution guesses the index set $I$. The substitution automaton that accomplishes this product sparsification can be found in Appendix B.1. Since the index set $I\subseteq[D_{2}]$ is unknown, the automaton non-deterministically selects which $\hat{Q}_{ij}$ polynomials will be treated as non-commutative. Given the structured nature of the polynomial $\hat{f}$, we can identify the boundary of each $\hat{Q}_{ij}$, ensuring that no additional spurious monomials are generated.
@@
$g_{J}=\sum_{i\in[s]}\left(\prod_{j\in\overline{J}}\hat{Q}_{ij}\right)\left(\prod_{j\in J}\hat{Q}_{ij}\right)\times\chi_{J}.$
@@
$\hat{f}^{{}^{\prime}}_{1}=\sum_{J\subseteq[D_{2}],|J|=s-1}g_{J}\times\chi_{J}.$
Moreover, $\hat{f}_{1}\neq 0$ if and only if $\hat{f}^{{}^{\prime}}_{1}=0$.
It’s evident that for distinct guesses $J$ and $J^{\prime}$ where $J\neq J^{\prime}$, the monomials of $g_{J}$ and $g_{J}^{\prime}$ will not mix, since the sub-monomials $\chi_{J}$ and $\chi_{J^{\prime}}$ are distinct. By Lemma 4.2.1, there exists index set $J\subseteq[D_{2}]$ with size at most $s-1$, such that $g_{J}\neq 0$ implying $\hat{f}_{1}^{\prime}\neq 0$.
@@
Let us denote the new polynomial obtained after replacing the commuting variables by scalars in $\hat{f}^{\prime}$ by $\hat{f}_{2}$.
@@ 4.3 Step 3: Commutative Transformation of $\hat{f}_{2}$
In this final step, we describe how to transform $\hat{f}_{2}$ into a commutative polynomial while preserving its non-zeroness. Note that $\hat{f}_{2}$ is a polynomial over $\mathbb{F}[\zeta\cup\chi]\langle\xi\rangle$. If we treat $\hat{f}_{2}$ as commutative by considering the n.c. variables $\xi$ as commutative, the exponents of the variable $\xi_{i}$ (for $i\in[s]$) from different n.c. $\hat{Q}_{ij}$ polynomials will be summed (or mixed). This mixing makes it impossible to guarantee that the resulting polynomial remains non-zero.
However, we can carefully convert $\hat{f}_{2}$ into a commutative polynomial while preserving its non-zeroness.
This is stated in the following lemma and the proof can be found in Appendix C. The substitution automaton for this step can be found in Appendix C.1.
Lemma 4.3.1.
Let $g=\sum_{i\in[s]}\beta_{i}\big{(}\prod_{j\in[s]}\hat{Q}_{ij})$, where $\beta_{i}\in\mathbb{F}$ and each $\hat{Q}_{ij}$ is an $s$-ordered power-sum polynomial over $\xi=\{\xi_{1},\xi_{2},\ldots,\xi_{s}\}$ of degree $D$. This can be expressed as: $g=\sum_{m}\alpha_{m}m$, where $m=\big{(}\prod_{j\in[s]}\xi_{1}^{i_{j1}}\xi_{2}^{i_{j2}}\ldots\xi_{s}^{i_{js}})$. The n.c. polynomial $g$ can be transformed into a commutative polynomial $g^{(c)}$ while preserving its non-zeroness.
@@
Similar to Lemma 4.2.1, Lemma 4.3.1 is more general because its proof relies only on the fact that each $\hat{Q}_{ij}$ polynomial is an $s$-ordered power-sum polynomial, and the number of products is bounded by some polynomial function in $s$. This allows us to generalize the lemma to sum of products of arbitrary $k$-ordered power-sum polynomials where $k\in\mathbb{N}$, with a proof that is analogous to that of the original lemma. We will utilize this generalization to transform polynomials computed by higher-depth $+$-regular circuits.

### v2

@@ 4.1.2 Non-zeroness of $\hat{f}$
Let $Z=\{z_{1},\ldots,z_{n}\}$ be the set of new commuting variables. Let $g\in\mathbb{F}\mathopen{\langle}X\mathclose{\rangle}$ be a polynomial of degree $D$ computed by a $\Sigma\Pi^{*}\Sigma$ circuit of size $s$. Then $g$ can be expressed as $g=\sum_{i\in[s]}\prod_{j\in[D]}L_{ij}$, where $L_{ij}$ are homogeneous linear forms. Let $P_{i}=\prod_{j\in[D]}L_{ij}$, $i\in[s]$. We have $g=\sum_{i\in[s]}P_{i}$. For $I\subseteq[D]$ with size at most $s-1$, define $P_{i,I}$ as the polynomial obtained from $P_{i}$ by treating linear forms indexed by $I$ as noncommuting and the rest of the linear forms as commuting. We replace each n.c. variable $x_{i}$ appearing in $[D]\setminus I$ by a new commuting variable $z_{i}$.
The number of n.c. linear forms appearing in $P_{i,I}\in\mathbb{F}[Z]\mathopen{\langle}X\mathclose{\rangle}$ is bounded by $|I|<s$. This is because the linear forms that appear with indices other than those in $I$ are treated as commutative. Consequently, the number of non-commutative linear forms in $P_{i,I}\in\mathbb{F}[Z]\langle X\rangle$ is bounded by $|I|<s$. We refer to this as the n.c. degree of the polynomial $P_{i,I}$. Since this degree is small, $P_{i,I}$ can be converted into a commutative polynomial while preserving its non-zeroness. Let $P^{(c)}_{i,I}$ denote the commutative polynomial obtained from $P_{i,I}$ and define $g_{I}=\sum_{i\in[s]}P^{(c)}_{i,I}$. To keep all guesses of the set $I$ distinct, additional commutative variables $\xi=\{\xi_{1},\xi_{2},\cdots,\xi_{k+1}\}$ are introduced in [AJMR19]. The transformed commutative polynomial obtained in [AJMR19] is given by:
@@
Let $f=\sum_{i\in[s]}\prod_{j\in[D_{2}]}Q_{ij}$ be a n.c. polynomial over $X=\{x_{1},\cdots,x_{n}\}$, computed by a $\Sigma\Pi^{*}\Sigma\Pi^{*}\Sigma$ circuit of size $s$. Let $D_{1}\geq s-1$ denote the degree of the polynomial $Q_{ij}$, $i\in[s],j\in[D_{2}]$. Let $\hat{f}\in\mathbb{F}[Y\sqcup Z]\mathopen{\langle}\xi\mathclose{\rangle}$ be defined as $\hat{f}=\sum_{i\in[s]}\prod_{j\in[D_{2}]}\hat{Q}_{ij},$ where $\hat{Q}_{ij}$ are as defined in Equation 6. Then, if $f\neq 0$ then $\hat{f}\neq 0$.
We will need a generalization of this Lemma 4.1.1 for polynomials computed by larger depth +-regular circuits. The generalization is stated in Lemma 5.1.1.
The resulting n.c. polynomial $\hat{f}$ (from Lemma 4.1.1) still has an exponential degree in $\xi$ variables, but each $\hat{Q}_{ij}$ is structured as $s$-ordered power-sum polynomials. Importantly, $\hat{f}$ does not contain any monomials from the spurious polynomial $F=\sum_{\begin{subarray}{c}m\in Mon(f)\end{subarray}}F_{m}$.
@@ 4.1.3 Small degree case
When $D_{1}<s-1$, we can use a small substitution automaton of size $c$, bounded by $s$, to substitute fresh n.c. double-indexed variables at each position within each $Q_{ij}$. Let $Z=\{z_{\ell k}\mid\ell\in[c]\text{ and }k\in[n]\}$ be the new set of n.c. variables. The automaton replaces the variable $x_{k}$ in the $\ell$-th position of a monomial of $Q_{ij}$ with $z_{\ell k}$ (this is a standard set-multilinear conversion of non-commutative polynomial $Q_{ij}$). To avoid introducing new notation for these transformed polynomials, we refer to the resulting polynomial by $\hat{Q}_{ij}$, which remains n.c. over $Z$. Unlike the high-degree case, in the small-degree case, the automaton can identify the boundary between $Q_{ij}$ and $Q_{i,j+1}$, ensuring that no spurious monomials are produced, that is, $F=0$.
@@
However, it is important to note that this transformation alone will not provide a black-box PIT, as we cannot guarantee the non-zeroness of the sum of products of these $\hat{Q}_{ij}$ polynomials. This is because if we simply treat all $\hat{Q}_{ij}$ as commutative, the variables across different $\hat{Q}_{ij}$ polynomials could mix, which may lead to cancellations. At this stage, the variables in $Z$ are still considered n.c. in the polynomial $\hat{f}$.
@@ 4.1.4 Non-zeroness of $f^{\prime}$
One of the goals of this transformation is to ensure that if $\xi_{s}$ is followed by $\xi_{1}$ in the transformed monomials (for all such occurrences of $\xi_{s}$ followed by $\xi_{1}$), there must be a transition from $Q_{i,j}$ to $Q_{i,j+1}$ for some $j\in[D_{2}]$. As noted, we cannot be sure of this. However, all those monomials where this transition occurs are captured in the structured part $\hat{f}_{1}$. Since there is no such structure in $F_{1}$, we cannot conclude anything about the monomials appearing in the spurious part $F_{1}$.
@@ 4.2 Step 2: Product Sparsification
In the second step of our transformation, we prove a general lemma that states if we have a sum of a small number of products of ordered power-sum polynomials, we can sparsify the product while preserving its non-zero property. Specifically, in each product term of the sum, we can treat only a small number of the ordered power-sum polynomials as non-commutative while treating the rest as commutative without affecting the non-zero nature of the polynomial. In particular, this step does not depend on the number of terms in each product.
We focus on the sparsification of the n.c. polynomial $\tilde{f}\in\mathbb{F}\mathopen{\langle}\xi\mathclose{\rangle}$, which was the output of Step (1). This transformation affects both the good part $\hat{f}_{1}$ and the spurious part $F_{1}$ of the polynomial $\tilde{f}\in\mathbb{F}\mathopen{\langle}\xi\mathclose{\rangle}$. We begin by analyzing the transformation of $\hat{f}_{1}$, which is defined as:
@@
where each $\hat{Q}_{ij}$ is an $s$-ordered power-sum polynomial in the n.c. variables $\xi=\{\xi_{1},\ldots,\xi_{s}\}$. Note that each $\hat{Q}_{ij}$ is a homogeneous and degree $D_{1}$ n.c. polynomial.
The key observation is that we can preserve the non-zeroness of $\hat{f}_{1}$ by retaining at most $s-1$ of the $s$-ordered power-sum polynomials $\hat{Q}_{ij}$ in each product $\prod_{j\in[D_{2}]}\hat{Q}_{ij}$ as non-commutative while treating the remaining ones as commutative. This is stated in the following lemma, which we refer to as the product sparsification lemma. This lemma generalizes Lemma 6.2 from [AJMR19]. However, unlike in [AJMR19], we are working with the product of non-commutative polynomials where the degree of individual factors can be greater than 1. If we simply treat them as commutative, as in [AJMR19], they may become zero. The proof of this lemma crucially relies on Claim 4.0.1, and the complete proof can be found in Appendix B. Unlike [AJMR19], one of the key distinctions in our setting is that the $\hat{Q}_{ij}$ polynomial can be non-homogeneous in general.
We will revisit the combination of $\hat{f}_{1}$ and $F_{1}$ in LABEL:{step1.1} to complete our analysis.
@@
where each $\hat{Q}_{ij}$ is an $s$-ordered power-sum polynomial of degree $D_{1}$ over $\xi=\{\xi_{1},\xi_{2},\ldots,\xi_{s}\}$. Then, there exists a subset $I\subseteq[D_{2}]$ with size at most $s-1$ such that if we treat the polynomials $\hat{Q}_{ij}$ for $j\in I$, as non-commutative and the others ($j\not\in I$) as commutative, then the polynomial $\hat{f}_{1}$ remains non-zero. Furthermore, each $\hat{Q}_{ij}$ polynomial may be non-homogeneous in general. Moreover, there is a small substitution automaton of size $O(s)$ that performs this transformation.
In other words, for each n.c. variable $\xi_{i}$ (where $i\in[s]$) in $\hat{f}_{1}$, there exists an $O(s)$-dimensional matrix—acting as a transition matrix of a substitution automaton of size $O(s)$. By evaluating $\hat{f}_{1}$ on these matrices, the polynomial is transformed into a product-sparsified polynomial while maintaining its non-zero property.
@@
We remark that the proof of Lemma 4.2.1 (see Appendix B) relies solely on the fact that each $\hat{Q}_{ij}$ polynomial in $\hat{f}_{1}$ is an ordered power-sum polynomial. In particular, the proof does not depend on the fact that each $\hat{Q}_{ij}$ is obtained from a $\Sigma\Pi^{*}\Sigma$ circuit. Instead, the proof relies on the following facts:
  1. 1.
    the number of summands is small,
  2. 2.
    the boundary of each ordered power-sum polynomial can be efficiently identified using a small automaton
This makes it irrelevant where the $\hat{Q}_{ij}$ polynomials originate from. As a result, we can apply this result whenever the given polynomial is represented as a sum of a small number of products of ordered power-sum polynomials (i.e., the number of summands is small). We will use this observation when working with higher-depth $+$-regular circuits.
Since we do not know the index set $I$, the substitution automaton guesses the index set $I$. The substitution automaton that accomplishes this product sparsification can be found in Appendix B.1. Since the index set $I\subseteq[D_{2}]$ is unknown, the automaton non-deterministically selects which $\hat{Q}_{ij}$ polynomials will be treated as non-commutative. Given the structured nature of the polynomial $\hat{f}$, we can identify the boundary of each $\hat{Q}_{ij}$, ensuring that no additional spurious monomials are generated.
@@
$g_{J}=\sum_{i\in[s]}\left(\prod_{j\in\overline{J}}\hat{Q}_{ij}\right)\left(\prod_{j\in J}\hat{Q}_{ij}\right)\times\chi_{J}.$ (9)
@@
$\hat{f}^{{}^{\prime}}_{1}=\sum_{J\subseteq[D_{2}],|J|=s-1}g_{J}.$
Moreover, $\hat{f}_{1}\neq 0$ if and only if $\hat{f}^{{}^{\prime}}_{1}\neq 0$.
It’s evident that for distinct guesses $J$ and $J^{\prime}$ where $J\neq J^{\prime}$, the monomials of $g_{J}$ and $g_{J}^{\prime}$ will not mix, since the sub-monomials $\chi_{J}$ and $\chi_{J^{\prime}}$ are distinct (see Equation 9). By Lemma 4.2.1, there exists index set $J\subseteq[D_{2}]$ with size at most $s-1$, such that $g_{J}\neq 0$ implying $\hat{f}_{1}^{\prime}\neq 0$.
@@
Let us denote the new polynomial obtained after replacing the commuting variables by scalars in $\hat{f_{1}}^{\prime}$ by $\hat{f}_{2}$.
@@ 4.3 Step 3: Commutative Transformation of $\hat{f}_{2}$
In this final step, we prove a general commutative transformation lemma, which states that if we have a non-commutative polynomial represented as a sum of products of a small number of ordered power-sum polynomials (i.e., the number of terms in each product is small), we can convert it into a commutative polynomial while preserving its non-zeroness property. In particular, this step does not depend on the number of summands. The key idea is to introduce a small number of new commutative variables to perform this transformation.
We now describe how to transform $\hat{f}_{2}$ into a commutative polynomial while preserving its non-zeroness. Note that $\hat{f}_{2}$ is a polynomial over $\mathbb{F}\langle\xi\rangle$. If we treat $\hat{f}_{2}$ as commutative by considering the n.c. variables $\xi$ as commutative, the exponents of the variable $\xi_{i}$ (for $i\in[s]$) from different n.c. $\hat{Q}_{ij}$ polynomials will be summed (or mixed). This mixing makes it impossible to guarantee that the resulting polynomial remains non-zero.
However, we can carefully convert $\hat{f}_{2}$ into a commutative polynomial while preserving its non-zeroness. This is stated in the following lemma and the proof can be found in Appendix C. In particular, there is a substitution automaton of size $O(s^{2})$ that carries out this commutative transformation and the substitution automaton is provided in Appendix C.1.
Lemma 4.3.1 ( Commutative Transformation Lemma).
Let $g=\sum_{i\in[s]}\beta_{i}\big{(}\prod_{j\in[s]}\hat{Q}_{ij})$, where $\beta_{i}\in\mathbb{F}$ and each $\hat{Q}_{ij}$ is an $s$-ordered power-sum polynomial over $\xi=\{\xi_{1},\xi_{2},\ldots,\xi_{s}\}$ of degree $D$. This can be expressed as: $g=\sum_{m}\alpha_{m}m$, where $m=\big{(}\prod_{j\in[s]}\xi_{1}^{i_{j1}}\xi_{2}^{i_{j2}}\ldots\xi_{s}^{i_{js}})$. The n.c. polynomial $g$ can be transformed into a commutative polynomial $g^{(c)}$ while preserving its non-zeroness. In particular, there exists a small substitution automaton of size $O(s^{2})$ that performs this commutative transformation.
In other words, for each n.c. variable $\xi_{i}$ (where $i\in[s]$), there exists an $O(s^{2})$-dimensional matrix—acting as a transition matrix of a substitution automaton. By evaluating $g$ on these matrices, the polynomial is transformed into a commutative polynomial $g^{(c)}$, while maintaining its non-zero property.
@@
We remark that Lemma 4.3.1 is more general. The proof (see Appendix C) depends only on the fact that the given non-commutative polynomial can be represented as a sum of products of a small number of ordered power-sum polynomials (i.e., the number of terms in each product is small). Crucially, the proof of this commutative transformation does not depend on the fact that the polynomials are derived from $+$-regular circuits or whether they are homogeneous. Instead, the proof relies on the following two facts:
  1. 1.
    The number of terms in each product is small, and
  2. 2.
    Each term in the product is an ordered power-sum polynomial so that boundaries can be identified efficiently using an automaton.
This makes the result applicable whenever the given non-commutative polynomial is represented as a sum of products of a small number of ordered power-sum polynomials. We will use this observation when working with higher-depth $+$-regular circuits.